# Geometry Aware Operator Transformer As An Efficient And Accurate Neural Surrogate For PDEs On Arbitrary Domains

### v3

@@ 1 Introduction
Partial Differential Equations (PDEs) are widely used to mathematically model very diverse natural and engineering systems PDEbook . Currently, numerical algorithms, such as the finite element and finite difference methods, are the preferred framework for simulating PDEs NAbook . However, these methods can be computationally very expensive, particularly for the so-called many-query problems such as uncertainty quantification (UQ), control, and inverse problems. Here, the solver must be called repeatedly, leading to prohibitive costs and providing the impetus for the design of fast and efficient surrogates for PDE solvers scimlrev2 .
To this end, ML/AI based algorithms are increasingly being explored as neural PDE surrogates. In particular, neural operators kovachki2023neural ; bartolucci2023representation , including those proposed in FNO ; GNO ; DeepONet ; CNO ; herde2024poseidon , which learn the PDE solution operator from data, are widely used azizzadenesheli2024neural . As many of these neural operators are restricted to PDEs on Cartesian (regular) grids, they cannot be directly applied to most engineering and industrial systems, which are set on domains with complex geometries, imposing a pressing need for neural operators for learning PDEs on arbitrary domains (point clouds).
In this context, a variety of options have recently been proposed, including domain masking for FNO and CNO CNO , replacing FFT in FNO with direct spectral evaluations lingsch2024beyond , augmenting FNO with learned diffeomorphisms geofno and mapping arbitrary point cloud data between the input domain and latent regular grids with learned encoders/decoders, while processing on the latent grid with FNO li2023geometryinformed or transformers UPT ; zhou2025textpde . Alternatives such as end-to-end message-passing based graph neural networks MGN ; MSMGN ; sanchez2018graph ; sanchez2020learning ; brandstetter2022message ; equer2023multi ; rigno or end-to-end transformer based models wu2024transolver ; luo2025transolver ; hao2023gnot ; wang2025cvit ; serrano2024aroma have also been proposed.
@@
A thorough comparison of the existing models, not just in terms of accuracy but also computational efficiency and scalability, is necessary to evaluate whether these models are yet capable enough to act as surrogates for highly successful finite element methods for engineering simulations NAbook . As one of the contributions in this paper, we performed such a comparison (see Sec.1 for details) to find evidence for an accuracy-efficiency trade-off, i.e., accurate and robust models, such as the message passing based RIGNO of rigno are not necessarily computationally efficient nor scalable in terms of training throughput and inference latency. On the other hand, more efficient models such as GINO li2023geometryinformed are not accurate enough (see the accompanying Radar Chart in Fig. 1). Given this observation, our main goal in this paper is to propose an algorithm to learn PDE solution operators on arbitrary domains that is accurate, computationally efficient, and can be seamlessly scaled to real-world industrial simulations.
To this end, we propose Geometry Aware Operator Transformer (GAOT, pronounced goat) as a neural surrogate for PDEs on arbitrary domains. While being based on the well-established encode-process-decode paradigm MGN , GAOT includes several novel features that are designed to ensure both computational efficiency and accuracy, such as
@@
  • A transformer processor that utilizes patching (as in ViTs dosovitskiy2021an ) for computational efficiency.
@@
  • The efficiency and scalability of GAOT is further showcased by it achieving state of the art (SOTA) performance on the large scale three-dimensional industrial benchmark of DrivAerNet++ dataset for automobile aerodynamics dvnet++ . We also test GAOT and demonstrate its superior performance to the GINO baseline on two further 3D industrial scale datasets, i.e., DrivaerML dataset for automobile aerodynamics and a NASA-CRM dataset for aerospace applications.
@@
The overall architecture of GAOT is depicted in Fig. 2. For simplicity of exposition, we start with the time-independent case, where given inputs $a(x_{j})$ on the point cloud $D_{\Delta}=\{x_{j}\}\subset D$, for $1\leq j\leq J$, GAOT provides an approximation to solution $u$ of the PDE (1) at any query point $x\in D$. At a high level, GAOT follows the encode-process-decode paradigm of MGN . In the first step, an encoder transforms the input on the underlying point cloud $D_{\Delta}$ to a latent point cloud $\mathcal{D}\subset\mathbb{R}^{d}$. The resulting spatial tokens are then processed by a processor module to learn useful representations and its output is remapped to the original domain $D$ via the decoder, which allows evaluation at any query point $x\in D$.
@@
As depicted in SM Fig. B.1, the latent domain $\mathcal{D}$ (to which the encoder maps) can be chosen in three different ways, namely i) a regular (structured) grid stencil, consisting of equispaced points on a Cartesian domain (see also Fig. 2) ii) randomly downsampling the underlying point cloud $D_{\Delta}$ or iii) a projected low-dimensional representation, where a high-dimensional domain is projected to a lower dimension (for instance using tri-plane embeddings in 3-D tripnet ) and a regular grid is used in the lower-dimensional domain. GAOT is a general framework where any of these latent point cloud choices can be employed for $\mathcal{D}$.
@@
Given input values $a(x_{j})$ on the underlying point cloud $D_{\Delta}$, the encoder aims to transform it into latent features $w_{e}(y)$ at any point $y\in\mathcal{D}$ on the latent point cloud. Using a graph-neural operator (GNO) encoder as in GINO li2023geometryinformed would lead to,
@@
The only geometric information in the afore-described encoder is provided by the coordinates of the underlying points. This alone does not convey the rich geometric information about the domain that can affect the solution of the underlying PDE (2). Hence, we need to embed further geometric information into the model. Deviating from the literature where geometric information is provided either by appending them as node and edge features on the underlying graphs MSMGN or by encoding a signed distance function li2023geometryinformed , we propose to use novel geometry embeddings to encode this information. To this end and as described in SM Sec. B.3, we can rely on local statistical embeddings for each point $y\in\mathcal{D}$ as all the neighboring points $x_{k}$ in $D_{\Delta}$ with $|y-x_{k}|\leq r_{m}$ have already been computed in the AGNO encoder. From these points, we can readily compute statistical descriptors such as i) number of neighbors $x_{k}\in D_{\Delta}$, in the ball $B_{r_{m}}(y)$, ii) the average distance $D_{{\rm avg}}=\frac{1}{n^{m}_{y}}\sum_{k=1}^{n_{y}^{m}}|y-x_{k}|$, iii) the variance of this distance $D_{{\rm var}}$, with respect to the average $D_{{\rm avg}}$, iv) the centroid offset vector $\Delta_{y}=\frac{1}{n^{m}_{y}}\sum_{k=1}^{n_{x}^{y}}(x_{k}-y)$ and v) a few principal component (PCA) features of the covariance matrix of $y-x_{k}$ to calculate the local shape anisotropy. These statistical descriptors, for each scale $m$ and each point $y\in\mathcal{D}$ are then concatenated into a single vector $z_{y}$, normalized across components to yield zero mean and unit variance and fed into an MLP to provide the embedding $g^{m}(y)$. Alternatively, geometry embedding using PointNet models Pnet can also be considered.
@@
The encoder provides a set of geometry aware tokens $w_{e}(y_{\ell})$, for all points $y_{\ell}\in\mathcal{D}$, with $1\leq\ell\leq L$, in the latent point cloud. These tokens are further transformed by a processor. As shown in Fig. 2, we choose a suitable transformer based processor. While postponing details on the processor architecture to SM Sec. B.4, we summarize our choices here. If the latent points $\{y_{\ell}\}$ lie on a regular grid (either through a structured stencil or a projected low-dimensional one), we use a patch-based vision transformer or ViT (dosovitskiy2021an and herde2024poseidon ; genCFD for PDE operator learning) for computational efficiency. The equispaced latent points are combined into patches and the tokens in each patch are flattened into a single token embedding which serves as the input for a multi-head attention block, followed by a feed forward block. RMS normalization is applied to the tokens before processing. Either sinusoidal absolute position embeddings or rotary relative position embeddings are used to encode token positions. If the latent points $y_{\ell}$ are randomly downsampled from the original point cloud, there is no obvious way to patch them together. Hence, a standard transformer vaswani2017attention , but with RMS normalization, can be used. Additionally, we employ multiple skip connections across transformer blocks (see Fig. 2). The transformer processor transforms the tokens $w_{e}(y_{\ell})$ into processed tokens, that we denote by $w_{p}(y_{\ell})$, for all $1\leq\ell\leq L$.
@@
To learn the solution operator $\mathcal{S}_{t}$ of the time-dependent PDE, we observe that the $\mathcal{S}_{t}$ can be used to update the solution forward in time, given the solution at any time point $u(t)$ by applying $u(t+\tau)=\mathcal{S}_{\tau}(u(t))$. Thus, for any time $t$, given the augmented input $a(t)=(c,u(t))$, with $c$ being the coefficient in the PDE (2), we need GAOT to output $u(t+\tau)$, for any $\tau\geq 0$. To this end, we retain the architecture of GAOT, as described for the time-independent case above, and simply add the current time $t$ and the lead-time $\tau$ as further inputs to the model. More precisely, the time-dependent version of GAOT is of the form $\widehat{\mathcal{S}}_{\theta}(x,t,\tau,a(t))$, where $a(t)$ takes values at points sampled in $D$. Following rigno , the map $\hat{S}_{\theta}$ can be used to update an approximate solution of PDE (2) in time by following a very general time-stepping strategy:
@@
Here, choosing the parameters $(\gamma,\delta)$ appropriately leads to different strategies for time stepping: $\gamma=0,\delta=1$ directly approximates the output of the solution operator at time $t+\tau$; $\gamma=1,\delta=1$ yields the residual of the solution at the later time, with respect to the solution at current time; $\gamma=1,\delta=\tau$ is equivalent to approximating the time-derivative of the solution. GAOT provides the flexibility to use any of these time-stepping strategies. We also use the all2all training strategy herde2024poseidon to leverage trajectory data for time-dependent PDEs.
@@
Datasets and Baselines. We start by testing GAOT of a challenging suite of 15 datasets for PDEs with input/output data on arbitrary point clouds in two space dimensions, see SM Secs D and G for a detailed description of the datasets and for visualizations, respectively. For time-independent PDEs, in addition to the elasticity benchmark of geofno , we consider two Poisson equation datasets: Poisson-Gauss, defined on random points in a square domain, and Poisson-C-Sines, a new dataset we propose, containing rich multiscale solutions on a circular domain. In addition, we propose 4 new datasets comprising compressible flows past objects, both airfoils as well as bluff bodies. These datasets have significant variation in domain geometry and flow conditions (Mach numbers ranging from subsonic to supersonic, varying angles of attack etc.) and are tailor-made for testing neural PDE surrogates on arbitrary domains in two space dimensions. For time-dependent PDEs, we test on the challenging datasets considered recently in rigno , composed of 8 operators corresponding to the compressible Euler (2), incompressible Navier-Stokes (4), and acoustic wave equations (2). These time-dependent operators include complex multiscale solutions with shocks and other sharp traveling waves which can interact, reflect and diffract making them hard to learn. We test GAOT on these datasets and compare them with several widely used neural operators for PDEs on arbitrary domains including those based on message passing (RIGNO rigno ), Fourier Layers (GINO li2023geometryinformed , GeoFNO geofno , FNO DSElingsch2024beyond ) and Transformers (Transolver wu2024transolver , UPT UPT and GNOT hao2023gnot ).
@@
How the training throughput of a model scales with increasing input and model size, is absolutely crucial for evaluating whether it can be used to process large-scale datasets (input scalability) or whether it can serve as a backbone of foundation models (model scalability) which require large model sizes herde2024poseidon . To evaluate the scalability of different models, we plot how the training throughput changes as input size and model size (Fig. 3 (a, b)) are increased to find that GAOT scales much more favorably than the baselines with respect to both input and model size. In fact, models like Transolver and GNOT scale very poorly, making it impossible for us to train them for the large-scale time-dependent datasets with all2all training, which requires handling large volumes of data for large input sizes. Hence, we omit them in the accuracy results for time-dependent datasets in Tab. 1. The results for both accuracy and efficiency across a range of metrics for GAOT, RIGNO, GINO and Transolver are summarized in SM Tab. E.4 and visualized in the Radar chart Fig. 1. This demonstrates that GAOT ensures both accuracy (robustness) and computational efficiency (scalability) at the same time, while being the best model performing model on both sets of metrics.
@@
Given the high accuracy and excellent computational efficiency and scalability of GAOT, we showcase its abilities further on three challenging three-dimensional large-scale benchmarks for industrial simulations. We start with the DrivAerNet++ dataset of dvnet++ . In this benchmark, the data consists of high-fidelity CFD simulations across 88K different car shapes which span the entire range of conventional car design. The underlying task is to learn steady-state surface fields (See Fig. 4) such as the pressure and wall shear stress, given the input car shape and flow conditions. The data has approximately 500500K points per shape, making the overall training extremely compute intensive. Thus, only scalable models can currently process this learning task. We test GAOT on this challenging 3D benchmark and report the RMSE and MAE test errors for the pressure and wall shear stress in Tab. 3. Compared to baselines results taken from the leaderboard of the DrivAerNet++ challenge tripnet , we see that GAOT significantly improves on the state-of-the-art (see also Fig. 4). This improvement is most visible in the MAE for wall shear stress where GAOT is ca. $30\%$ more accurate than the second-best model (TripNet), which currently sits atop the leaderboard for wall shear stress predictions. We recall that GAOT’s decoder endows it with neural field properties. We showcase it for the DrivAerNet++ dataset by training GAOT on a randomly selected set of less than $10\%$ of the total input points (per batch) and then testing on the original car surface point cloud by querying the desired points through GAOT’s decoder. Although not as accurate as training GAOT with full input, we observe from Tab. 3 that this neural field version of GAOT has comparable accuracy to some of the baselines which have been trained with 1010x more input points, further demonstrating the flexibility and accuracy of GAOT. Further assessments of the neural field property of GAOT are provided in SM E.8.
@@
Next, we consider the very challenging DrivAerML dataset of ashton2024drivaer (see Fig. 5), where the learning task is exactly the same as in DrivAernet++, i.e., predicting surface fields on the car, given its shape as the input. However, unlike DrivAernet++ which was based on coarse RANS simulations, DrivAerML’s ground truth is based on highly accurate LES simulations. This enables the incorporation of much more fine-scale physical effects in this dataset. The learning problem becomes harder, not just in terms of the challenging underlying physics, but also the fact that the number of points on the car surface is now 9M, instead of 500K for DrivAernet++. Thus, only highly scalable models can deal with this extreme resolution. Consequently, we are only able to test GAOT and GINO for this dataset and report the results in Table 4 to observe that GAOT is significantly (almost twice on wall shear-stress) as accurate as GINO on this dataset.
@@
Finally, we consider an industrial-scale dataset, recently proposed in bekemeyer2025introduction , where the learning task (see Fig. 5), is to predict surface pressure and the skin friction coefficient, given the shape of a full aircraft. The ground truth is generated with RANS simulations using a Spalart-Allmaras turbulence model, and the results with GAOT and GINO are reported in Table 4, showing that GAOT significantly outperforms GINO on this large-scale industrial dataset.
@@
We highlight GAOT’s flexibility with respect to the point distributions that it can handle by testing it on PDEs with regular grid inputs, as suggested in rigno . To this end, we considered 7 additional datasets and present the test errors in SM Sec. E.5. to find that GAOT is highly accurate even on regular grids and is either more accurate or comparable to the highly expressive GNN-based RIGNO, while being more accurate than widely used neural operators such as FNO and CNO. A key requirement in operator learning kovachki2023neural ; bartolucci2023representation is the ability of the model to generalize (at test time) to input and output resolutions that are different from the training resolution. As GAOT can be readily evaluated at any query point, we showcase this aspect of GAOT in SM Sec. E.7. by plotting the test errors for a sequence of resolutions, different from the training resolution, for the Poisson-Gauss benchmark, to find that GAOT generalizes very well in both the sub- and super-resolution settings, even to grids with 1010x more input points than the training resolution. Another test of the generalizability of a model is its ability to perform well out-of-distribution, either zero-shot or when it is fine-tuned with a few in-distribution samples for the new learning task. To test this aspect, we consider the datasets for compressible flow past bluff bodies and train a GAOT model on a set of bluff body shapes and then test it on shapes that were not in the training set. Then, the model is fine-tuned with a few task-specific samples and the results are shown in Fig. 3 (c). We observe that our model performs very well in a few-shot transfer learning scenario, with the fine-tuned model providing an almost order of magnitude gain in accuracy over the model, trained from scratch. Finally, in SM Sec. E.6, we demonstrate that GAOT scales with both model and dataset size, with scaling with respect to dataset size, also illustrated in Fig. 3 (c)

### v4

@@ 1 Introduction
Partial Differential Equations (PDEs) are widely used to mathematically model very diverse natural and engineering systems Evans (2022). Currently, numerical algorithms, such as the finite element and finite difference methods, are the preferred framework for simulating PDEs Quarteroni and Valli (1994). However, these methods can be computationally very expensive, particularly for the so-called many-query problems such as uncertainty quantification (UQ), control, and inverse problems. Here, the solver must be called repeatedly, leading to prohibitive costs and providing the impetus for the design of fast and efficient surrogates for PDE solvers Mishra and Townsend (2024).
To this end, ML/AI based algorithms are increasingly being explored as neural PDE surrogates. In particular, neural operators Kovachki et al. (2023); Bartolucci et al. (2023), including those proposed in Li et al. (2020a, b); Lu et al. (2021); Raonic et al. (2023); Herde et al. (2024), which learn the PDE solution operator from data, are widely used Azizzadenesheli et al. (2024). As many of these neural operators are restricted to PDEs on Cartesian (regular) grids, they cannot be directly applied to most engineering and industrial systems, which are set on domains with complex geometries, imposing a pressing need for neural operators for learning PDEs on arbitrary domains (point clouds).
In this context, a variety of options have recently been proposed, including domain masking for FNO and CNO Raonic et al. (2023), replacing FFT in FNO with direct spectral evaluations Lingsch et al. (2024), augmenting FNO with learned diffeomorphisms Li et al. (2023a) and mapping arbitrary point cloud data between the input domain and latent regular grids with learned encoders/decoders, while processing on the latent grid with FNO Li et al. (2023b) or transformers Alkin et al. (2024); Zhou et al. (2025). Alternatives such as end-to-end message-passing based graph neural networks Pfaff et al. (2021); Fortunato et al. (2022); Sanchez-Gonzalez et al. (2018, 2020); Brandstetter et al. (2022); Equer et al. (2023); Mousavi et al. (2025) or end-to-end transformer based models Wu et al. (2024); Luo et al. (2025); Hao et al. (2023); Wang et al. (2025); Serrano et al. (2024) have also been proposed.
@@
A thorough comparison of the existing models, not just in terms of accuracy but also computational efficiency and scalability, is necessary to evaluate whether these models are yet capable enough to act as surrogates for highly successful finite element methods for engineering simulations Quarteroni and Valli (1994). As one of the contributions in this paper, we performed such a comparison (see Sec.1 for details) to find evidence for an accuracy-efficiency trade-off, i.e., accurate and robust models, such as the message passing based RIGNO of Mousavi et al. (2025) are not necessarily computationally efficient nor scalable in terms of training throughput and inference latency. On the other hand, more efficient models such as GINO Li et al. (2023b) are not accurate enough (see the accompanying Radar Chart in Fig. 1). Given this observation, our main goal in this paper is to propose an algorithm to learn PDE solution operators on arbitrary domains that is accurate, computationally efficient, and can be seamlessly scaled to real-world industrial simulations.
To this end, we propose Geometry Aware Operator Transformer (GAOT, pronounced goat) as a neural surrogate for PDEs on arbitrary domains. While being based on the well-established encode-process-decode paradigm Pfaff et al. (2021), GAOT includes several novel features that are designed to ensure both computational efficiency and accuracy, such as
@@
  • A transformer processor that utilizes patching (as in ViTs Dosovitskiy et al. (2021)) for computational efficiency.
@@
  • The efficiency and scalability of GAOT is further showcased by it achieving state of the art (SOTA) performance on the large scale three-dimensional industrial benchmark of DrivAerNet++ dataset for automobile aerodynamics Elrefaie et al. (2025a). We also test GAOT and demonstrate its superior performance to the GINO baseline on two further 3D industrial scale datasets, i.e., DrivaerML dataset for automobile aerodynamics and a NASA-CRM dataset for aerospace applications.
@@
The overall architecture of GAOT is depicted in Fig. 2. For simplicity of exposition, we start with the time-independent case, where given inputs $a(x_{j})$ on the point cloud $D_{\Delta}=\{x_{j}\}\subset D$, for $1\leq j\leq J$, GAOT provides an approximation to solution $u$ of the PDE (1) at any query point $x\in D$. At a high level, GAOT follows the encode-process-decode paradigm of Pfaff et al. (2021). In the first step, an encoder transforms the input on the underlying point cloud $D_{\Delta}$ to a latent point cloud $\mathcal{D}\subset\mathbb{R}^{d}$. The resulting spatial tokens are then processed by a processor module to learn useful representations and its output is remapped to the original domain $D$ via the decoder, which allows evaluation at any query point $x\in D$.
@@
As depicted in SM Fig. B.1, the latent domain $\mathcal{D}$ (to which the encoder maps) can be chosen in three different ways, namely i) a regular (structured) grid stencil, consisting of equispaced points on a Cartesian domain (see also Fig. 2) ii) randomly downsampling the underlying point cloud $D_{\Delta}$ or iii) a projected low-dimensional representation, where a high-dimensional domain is projected to a lower dimension (for instance using tri-plane embeddings in 3-D Chen et al. (2025)) and a regular grid is used in the lower-dimensional domain. GAOT is a general framework where any of these latent point cloud choices can be employed for $\mathcal{D}$.
@@
Given input values $a(x_{j})$ on the underlying point cloud $D_{\Delta}$, the encoder aims to transform it into latent features $w_{e}(y)$ at any point $y\in\mathcal{D}$ on the latent point cloud. Using a graph-neural operator (GNO) encoder as in GINO Li et al. (2023b) would lead to,
@@
The only geometric information in the afore-described encoder is provided by the coordinates of the underlying points. This alone does not convey the rich geometric information about the domain that can affect the solution of the underlying PDE (2). Hence, we need to embed further geometric information into the model. Deviating from the literature where geometric information is provided either by appending them as node and edge features on the underlying graphs Fortunato et al. (2022) or by encoding a signed distance function Li et al. (2023b), we propose to use novel geometry embeddings to encode this information. To this end and as described in SM Sec. B.3, we can rely on local statistical embeddings for each point $y\in\mathcal{D}$ as all the neighboring points $x_{k}$ in $D_{\Delta}$ with $|y-x_{k}|\leq r_{m}$ have already been computed in the AGNO encoder. From these points, we can readily compute statistical descriptors such as i) number of neighbors $x_{k}\in D_{\Delta}$, in the ball $B_{r_{m}}(y)$, ii) the average distance $D_{{\rm avg}}=\frac{1}{n^{m}_{y}}\sum_{k=1}^{n_{y}^{m}}|y-x_{k}|$, iii) the variance of this distance $D_{{\rm var}}$, with respect to the average $D_{{\rm avg}}$, iv) the centroid offset vector $\Delta_{y}=\frac{1}{n^{m}_{y}}\sum_{k=1}^{n_{x}^{y}}(x_{k}-y)$ and v) a few principal component (PCA) features of the covariance matrix of $y-x_{k}$ to calculate the local shape anisotropy. These statistical descriptors, for each scale $m$ and each point $y\in\mathcal{D}$ are then concatenated into a single vector $z_{y}$, normalized across components to yield zero mean and unit variance and fed into an MLP to provide the embedding $g^{m}(y)$. Alternatively, geometry embedding using PointNet models Qi et al. (2017b) can also be considered.
@@
The encoder provides a set of geometry aware tokens $w_{e}(y_{\ell})$, for all points $y_{\ell}\in\mathcal{D}$, with $1\leq\ell\leq L$, in the latent point cloud. These tokens are further transformed by a processor. As shown in Fig. 2, we choose a suitable transformer based processor. While postponing details on the processor architecture to SM Sec. B.4, we summarize our choices here. If the latent points $\{y_{\ell}\}$ lie on a regular grid (either through a structured stencil or a projected low-dimensional one), we use a patch-based vision transformer or ViT (Dosovitskiy et al. (2021) and Herde et al. (2024); Molinaro et al. (2025) for PDE operator learning) for computational efficiency. The equispaced latent points are combined into patches and the tokens in each patch are flattened into a single token embedding which serves as the input for a multi-head attention block, followed by a feed forward block. RMS normalization is applied to the tokens before processing. Either sinusoidal absolute position embeddings or rotary relative position embeddings are used to encode token positions. If the latent points $y_{\ell}$ are randomly downsampled from the original point cloud, there is no obvious way to patch them together. Hence, a standard transformer Vaswani (2017), but with RMS normalization, can be used. Additionally, we employ multiple skip connections across transformer blocks (see Fig. 2). The transformer processor transforms the tokens $w_{e}(y_{\ell})$ into processed tokens, that we denote by $w_{p}(y_{\ell})$, for all $1\leq\ell\leq L$.
@@
To learn the solution operator $\mathcal{S}_{t}$ of the time-dependent PDE, we observe that the $\mathcal{S}_{t}$ can be used to update the solution forward in time, given the solution at any time point $u(t)$ by applying $u(t+\tau)=\mathcal{S}_{\tau}(u(t))$. Thus, for any time $t$, given the augmented input $a(t)=(c,u(t))$, with $c$ being the coefficient in the PDE (2), we need GAOT to output $u(t+\tau)$, for any $\tau\geq 0$. To this end, we retain the architecture of GAOT, as described for the time-independent case above, and simply add the current time $t$ and the lead-time $\tau$ as further inputs to the model. More precisely, the time-dependent version of GAOT is of the form $\widehat{\mathcal{S}}_{\theta}(x,t,\tau,a(t))$, where $a(t)$ takes values at points sampled in $D$. Following Mousavi et al. (2025), the map $\hat{S}_{\theta}$ can be used to update an approximate solution of PDE (2) in time by following a very general time-stepping strategy:
@@
Here, choosing the parameters $(\gamma,\delta)$ appropriately leads to different strategies for time stepping: $\gamma=0,\delta=1$ directly approximates the output of the solution operator at time $t+\tau$; $\gamma=1,\delta=1$ yields the residual of the solution at the later time, with respect to the solution at current time; $\gamma=1,\delta=\tau$ is equivalent to approximating the time-derivative of the solution. GAOT provides the flexibility to use any of these time-stepping strategies. We also use the all2all training strategy Herde et al. (2024) to leverage trajectory data for time-dependent PDEs.
@@
Datasets and Baselines. We start by testing GAOT of a challenging suite of 15 datasets for PDEs with input/output data on arbitrary point clouds in two space dimensions, see SM Secs D and G for a detailed description of the datasets and for visualizations, respectively. For time-independent PDEs, in addition to the elasticity benchmark of Li et al. (2023a), we consider two Poisson equation datasets: Poisson-Gauss, defined on random points in a square domain, and Poisson-C-Sines, a new dataset we propose, containing rich multiscale solutions on a circular domain. In addition, we propose 4 new datasets comprising compressible flows past objects, both airfoils as well as bluff bodies. These datasets have significant variation in domain geometry and flow conditions (Mach numbers ranging from subsonic to supersonic, varying angles of attack etc.) and are tailor-made for testing neural PDE surrogates on arbitrary domains in two space dimensions. For time-dependent PDEs, we test on the challenging datasets considered recently in Mousavi et al. (2025), composed of 8 operators corresponding to the compressible Euler (2), incompressible Navier-Stokes (4), and acoustic wave equations (2). These time-dependent operators include complex multiscale solutions with shocks and other sharp traveling waves which can interact, reflect and diffract making them hard to learn. We test GAOT on these datasets and compare them with several widely used neural operators for PDEs on arbitrary domains including those based on message passing (RIGNO Mousavi et al. (2025)), Fourier Layers (GINO Li et al. (2023b), GeoFNO Li et al. (2023a), FNO DSELingsch et al. (2024)) and Transformers (Transolver Wu et al. (2024), UPT Alkin et al. (2024) and GNOT Hao et al. (2023)).
@@
How the training throughput of a model scales with increasing input and model size, is absolutely crucial for evaluating whether it can be used to process large-scale datasets (input scalability) or whether it can serve as a backbone of foundation models (model scalability) which require large model sizes Herde et al. (2024). To evaluate the scalability of different models, we plot how the training throughput changes as input size and model size (Fig. 3 (a, b)) are increased to find that GAOT scales much more favorably than the baselines with respect to both input and model size. In fact, models like Transolver and GNOT scale very poorly, making it impossible for us to train them for the large-scale time-dependent datasets with all2all training, which requires handling large volumes of data for large input sizes. Hence, we omit them in the accuracy results for time-dependent datasets in Tab. 1. The results for both accuracy and efficiency across a range of metrics for GAOT, RIGNO, GINO and Transolver are summarized in SM Tab. E.4 and visualized in the Radar chart Fig. 1. This demonstrates that GAOT ensures both accuracy (robustness) and computational efficiency (scalability) at the same time, while being the best model performing model on both sets of metrics.
@@
Given the high accuracy and excellent computational efficiency and scalability of GAOT, we showcase its abilities further on three challenging three-dimensional large-scale benchmarks for industrial simulations. We start with the DrivAerNet++ dataset of Elrefaie et al. (2025a). In this benchmark, the data consists of high-fidelity CFD simulations across 88K different car shapes which span the entire range of conventional car design. The underlying task is to learn steady-state surface fields (See Fig. 4) such as the pressure and wall shear stress, given the input car shape and flow conditions. The data has approximately 500500K points per shape, making the overall training extremely compute intensive. Thus, only scalable models can currently process this learning task. We test GAOT on this challenging 3D benchmark and report the RMSE and MAE test errors for the pressure and wall shear stress in Tab. 3. Compared to baselines results taken from the leaderboard of the DrivAerNet++ challenge Chen et al. (2025), we see that GAOT significantly improves on the state-of-the-art (see also Fig. 4). This improvement is most visible in the MAE for wall shear stress where GAOT is ca. $30\%$ more accurate than the second-best model (TripNet), which currently sits atop the leaderboard for wall shear stress predictions. We recall that GAOT’s decoder endows it with neural field properties. We showcase it for the DrivAerNet++ dataset by training GAOT on a randomly selected set of less than $10\%$ of the total input points (per batch) and then testing on the original car surface point cloud by querying the desired points through GAOT’s decoder. Although not as accurate as training GAOT with full input, we observe from Tab. 3 that this neural field version of GAOT has comparable accuracy to some of the baselines which have been trained with 1010x more input points, further demonstrating the flexibility and accuracy of GAOT. Further assessments of the neural field property of GAOT are provided in SM E.8. A natural follow-up question in this regard is whether training at full (high) resolution provides a significant advantage over training at low resolution and inferring at high resolution ? Recent works such as Elrefaie et al. (2025b) have investigated this question in the context of the DrivAerNet++ dataset. We explore this issue in SM E.9 and demonstrate that GOAT trained at full resolution is significantly (almost twice) more accurate than competing models trained at low resolution and inferred at full resolution, see SM Table E.8.
@@
Next, we consider the very challenging DrivAerML dataset of Ashton et al. (2024) (see Fig. 5), where the learning task is exactly the same as in DrivAernet++, i.e., predicting surface fields on the car, given its shape as the input. However, unlike DrivAernet++ which was based on coarse RANS simulations, DrivAerML’s ground truth is based on highly accurate LES simulations. This enables the incorporation of much more fine-scale physical effects in this dataset. The learning problem becomes harder, not just in terms of the challenging underlying physics, but also the fact that the number of points on the car surface is now 9M, instead of 500K for DrivAernet++. Thus, only highly scalable models can deal with this extreme resolution. Consequently, we are only able to test GAOT and GINO for this dataset and report the results in Table 4 to observe that GAOT is significantly (almost twice on wall shear-stress) as accurate as GINO on this dataset.
@@
Finally, we consider an industrial-scale dataset, recently proposed in Bekemeyer et al. (2025), where the learning task (see Fig. 5), is to predict surface pressure and the skin friction coefficient, given the shape of a full aircraft. The ground truth is generated with RANS simulations using a Spalart-Allmaras turbulence model, and the results with GAOT and GINO are reported in Table 4, showing that GAOT significantly outperforms GINO on this large-scale industrial dataset.
@@
We highlight GAOT’s flexibility with respect to the point distributions that it can handle by testing it on PDEs with regular grid inputs, as suggested in Mousavi et al. (2025). To this end, we considered 7 additional datasets and present the test errors in SM Sec. E.5. to find that GAOT is highly accurate even on regular grids and is either more accurate or comparable to the highly expressive GNN-based RIGNO, while being more accurate than widely used neural operators such as FNO and CNO. A key requirement in operator learning Kovachki et al. (2023); Bartolucci et al. (2023) is the ability of the model to generalize (at test time) to input and output resolutions that are different from the training resolution. As GAOT can be readily evaluated at any query point, we showcase this aspect of GAOT in SM Sec. E.7. by plotting the test errors for a sequence of resolutions, different from the training resolution, for the Poisson-Gauss benchmark, to find that GAOT generalizes very well in both the sub- and super-resolution settings, even to grids with 1010x more input points than the training resolution. Another test of the generalizability of a model is its ability to perform well out-of-distribution, either zero-shot or when it is fine-tuned with a few in-distribution samples for the new learning task. To test this aspect, we consider the datasets for compressible flow past bluff bodies and train a GAOT model on a set of bluff body shapes and then test it on shapes that were not in the training set. Then, the model is fine-tuned with a few task-specific samples and the results are shown in Fig. 3 (c). We observe that our model performs very well in a few-shot transfer learning scenario, with the fine-tuned model providing an almost order of magnitude gain in accuracy over the model, trained from scratch. Finally, in SM Sec. E.6, we demonstrate that GAOT scales with both model and dataset size, with scaling with respect to dataset size, also illustrated in Fig. 3 (c)